x+y=20
30x+50y=880

Answer by mananth(16946)

1 x + 1 y = 20 .............1
30 x + 50 y = 880 .............2
Eliminate y
multiply (1)by -50
Multiply (2) by 1
-50 x -50 y = -1000
30 x + 50 y = 880
Add the two equations
-20 x = -120
/ -20
x = 6
plug value of x in (1)
1 x + 1 y = 20
6 + 1 y = 20
1 y = 20-6
1 y = 14
y = 14